# SocialEyes: Scaling mobile eye-tracking to multi-person social settings

### v2

@@ 4.1.3. Stimuli
  • Music Performance: American percussionist-composers, Allen Otte and John Lane performed the Canadian premiere of The Innocents – hour-long performance art piece centred on the issue of wrongful incarceration. The work might best be described as social justice advocacy through performance art. The hour-long composition consists of home-made instruments, spoken texts, and electronic soundscapes. Across 17 tableaux–many of which elide into one another–the audience is presented with dramatic soundscapes that explore issues related to the American criminal justice system, such as mistaken identity, injustice, truth, politics, psychology, and resilience. Some tableaux are too long, uncomfortable, confusing; others are direct, melodic, and in familiar styles.
  • Film: Otte and Lane’s The Innocents] is now the subject of an award-winning feature-length (1hr 20min) documentary film, directed by Wojciech Lorenc. The film showcases Otte and Lane’s compositional process and friendship, following them on the road as they tour the United States performing. On tour, we witness conversations with relevant stakeholders, such as lawyers, non-profit organisations, students, and exonerees. Woven throughout the film is the story of exoneree Anna Vasquez who served 13 years in prison for a crime that never occurred. We witness them tell parts of their story and ultimately their reaction to watching Otte and Lane’s live performance.
@@ 4.1.4. Study Design
The current utility test takes place within the context of a larger project examining the role of context information in altering audience physiological responses and social attitudes and behaviours. Specifically, counterbalancing across two different event days, we manipulated the order of the film vs. the performance [film-performance or performance-film]. Audience members could buy tickets for either event day and were aware that both the film and performance would occur; however, we never advertised the order, or the fact that we were manipulating order.
During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded eye measures, via web browser and webcam ((Saxena et al., 2022, 2023)). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our software framework–SocialEyes.
@@ 4.1.5. Procedure
To ensure proper connection to all eye-tracking devices, test recordings for a short length were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye-tracking devices (see Figure 4A).
@@ 4.2.1. Time Synchronisation
Overall, these data show relatively low offsets and drifts across all 30 devices and recording sessions. As is expected, mean offsets increase from the 1st to 2nd recording session for both groups (Fig. 5A; left panel), while drifts seem more so to be device-specific than recording session-specific. Regardless of the actual calculated offsets and drifts, our RANSAC approach allows us to accurately align the timestamps between each eye-tracking mobile device and our central camera.
@@ 4.2.2. Visualisation
Figure 6. Visualisations. A An example of feature matching from one frame of one participants’ worldview (left) onto the centralview (right). The recorded gaze corresponding to this frame is represented by the maroon and white circle. B Multi-person views generated with the visualisation module. Top: All participants’ egocentric worldviews and gazes (small white and black circles) are displayed around the perimeter. Using the computed homography matrices for each participant, we project each participant’s gaze onto the centralview (centre). In the centre panel, participants’ gazes are represented with uniquely coloured circles with white in the middle to aid visibility. Bottom: Same as top, with the centre grid cell displaying a heatmap of the 2D gaze density of all participants looking at the scene. Higher intensities in the heatmap (intensity increases from blue to red with red being the highest) represent a higher proportion of the participants looking at that location. C Same as centre panels in B, but resized to aid visibility.
@@
Raw egocentric gaze coordinates and homography-transformed gaze coordinates were projected onto each viewer’s worldview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 6) enable data validation and demonstrate accurate projective transformation using the SuperGlue features identified in the Homography module (see 3.2.6). The transforms handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and worldviews) reliably. However, further optimisations and investigations would be needed to identify edge cases in the homography estimation. Figure 6 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/multipersonET. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish audience-gaze on one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”); the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights the observation that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
We analyse the 2D distribution of gaze with the help of heatmaps (see Figures 7 and 8). Figure 7 shows the raw egocentric gaze data recorded from the eye-tracking glasses in the worldview coordinate space. This egocentric gaze data is relative to the wearer’s head movements and exhibits low variance in both horizontal and vertical directions with a strong central bias. This effect could be explained by the high coupling between eye and head movements commonly observed in real-world exploration (Einhäuser et al., 2007; Tatler, 2007).
@@ 4.2.3. Analysis
Further comparisons between individual heatmaps can be made with the help of metrics from our Analysis module (see 3.2.8). Pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 10; left column), nearly doubling in most cases. A 2D chart of seat positions (row, column) for each participant (see Figure 7 and 8 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., higher SIM values for smaller seat distances between participants in a pair and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P.
@@
Figure 10. Gaze entropy (top left), average gaze velocity (bottom left), pairwise heatmap similarity (top right) and pairwise heatmap correlation (bottom right) measures computed for each of the two sessions on both days. Coloured dots in the left plots represent individual participant-wise entropy and velocity measures. Coloured dots in the right plots represent SIM and CC for all unique participant-pairs. Black dots in all plots represent means and error bars represent confidence intervals for the respective mean values. Gaze entropy and velocity give a measure of individual gaze variability for each participant and heatmap similarity (SIM) and correlation ($C$) provide an estimate of similarities within participants’ gaze exploration for each session.
@@ 5. General Discussion
The proposed method conceptualised and tackled key challenges related to multi-person eye-tracking, yet a major external dependency of our framework is the choice of eye-tracking hardware. We acknowledge that the availability/development of required functionality in existing eye trackers and affordability are still substantial barriers. We picked Pupil Labs Neon devices in our utility test over market competitors because of their remote API control, minimal calibration requirements, and the availability of open-source gaze-processing components. However, in our initial tests we encountered an array of operational issues on the proprietary vendor application that severely hindered reliable recording from the devices. Solving these issues required hours of testing, troubleshooting or downscaling. For example, the Neon devices are also capable of recording device orientation and movement at 110 Hz using the onboard 9-DoF Inertial Measurement Unit (IMU), as well as stereo audio from on-board microphones; we had to omit these data streams to allow reliable recording of the gaze and video data streams, which were our priority. IMU failures would crash the mobile application, even though the eye data were intact, and audio recording could not be reliably remotely started through API calls, as it requires the android device to be unlocked. These issues point out the very incipient stage of multi-person eye-tracking hardware/software since these systems are mostly tested for single person use cases. Moving forward, the technology would benefit from increased heavy testing, at the scale we propose in this paper. Further, the availability of open-source implementations would allow higher customizability for application specific requirements and thereby contribute to its adoption in novel settings.
@@
Future objectives for our framework would be to further investigate and optimise gaze mapping accuracy in different settings. For example, identifying/handling edge failure cases in challenging environments and adding temporal information from sequential frames would help reduce prediction errors and noise in the projected gaze. The framework can also be extended to support additional social environments and hardware vendors. The current paper demonstrates gaze mapping in scenes where multiple people share a scene view; however, the application could be scaled up to social settings where individuals share different views of the same scene using Multi-view Multi-object detection and tracking (Taj and Cavallaro, 2010). Moreover, the performance bottleneck of homography estimation could be improved by using techniques like feature tracking with periodic detection updates to optimise computation which would pave the way for real-time use cases, such as gaze-contingent interaction and sonification (Hornof, 2014). Future development and applications can also incorporate multiple sensor types, such as physiological or IOT sensors, to record and analyse synchronous multimodal measures. We look forward to applying this framework to study attention and subjective experiences in immersive, social contexts, and plan to open-source the code to motivate collaboration and application of our framework in novel contexts.
@@ 6. Conclusion
7. Code Availability
All code will be made publicly available on GitHub upon acceptance/publication of the submitted manuscript.

### v3

@@ 4.1.3. Stimuli
  • Music Performance: The hour-long composition consisted of home-made instruments, spoken texts, and electronic soundscapes to present the audience with dramatic soundscapes that explore issues related to the American criminal justice system. Of the 17 total tableaux included in the performance, some were too long, uncomfortable, confusing; others were direct, melodic, and in familiar styles.
  • Film:The film showcases the compositional process behind the performance, engaging with exonerees, whose story and reaction to the performance is also captured.
@@ 4.1.4. Study Design
The current utility test takes place within the context of a larger project examining the role of context information in altering audience physiological responses and social attitudes and behaviours.During the event, in addition to wearing eye-tracking glasses, participants also wore a smart watch to monitor cardiac activity and completed a series of surveys. Parallel to this in-person event, we also presented a livestream online. We recorded eye measures, via web browser and webcam ((Saxena et al., 2022, 2023)). We were interested in the effect social co-presence might have on our measures of interest. Detailed analyses related to our cognitive neuroscientific and social psychological questions will be reported elsewhere. In the current study, we focus on validating the utility of our software framework–SocialEyes.
@@ 4.1.5. Procedure
Before the event, ADB connection was established between the server and each smartphone device. A forced synchronisation to the assigned NTP server on the smartphones was also done. Participants were seated evenly across rows two to four in the audience, hereby referred to as participant-rows one to three in this paper. After taking their seat, participants were instructed on how to put the glasses on. Trained research assistants completed a quick calibration procedure. Note that explicit eye-egoview calibration was not needed for the eye-tracking hardware (see 4.1.1). The calibration procedure mentioned here simply performs an individual offset correction using a single fixation target, which can account for physiological factors, such as the Kappa angle between the optical and visual axis of the eye. Participants retained their seating positions over the entire study duration.
To ensure proper connection to all eye-tracking devices, test recordings for a short length (1-5 minutes) were made before the actual recording session was started. We started the recording session before the event host came on stage. After the start of the event, no interruptions were allowed and all interactions with the recording devices were done remotely using the GlassesRecord module TUI to trigger recordings and monitor the eye-tracking devices (see GlassesRecord in 4.1.2).
@@ 4.2.1. Time Synchronisation
Overall, these data show relatively low offsets and drifts across all 30 devices and recording sessions. As is expected, mean offsets increase from the 1st to 2nd recording session for both groups (Figure 5A; left panel), while drifts seem more so to be device-specific than recording session-specific. Regardless of the actual calculated offsets and drifts, our RANSAC approach allows us to accurately align the timestamps between each eye-tracking mobile device and our central camera.
@@ 4.2.2. Visualisation
Figure 5. Visualisations. A An example of feature matching from one frame of one participants’ egoview (left) onto the centralview (right). The recorded gaze corresponding to this frame is represented by the maroon and white circle. B Multi-person views generated with the visualisation module. Top: All participants’ egoviews and gazes (small white and black circles) are displayed around the perimeter. Using the computed homography matrices for each participant, we project each participant’s gaze onto the centralview (centre). In the centre panel, participants’ gazes are represented with uniquely coloured circles with white in the middle to aid visibility. Bottom: Same as top, with the centre grid cell displaying a heatmap of the 2D gaze density of all participants looking at the scene. Higher intensities in the heatmap (intensity increases from blue to red with red being the highest) represent a higher proportion of the participants looking at that location. C Same as centre panels in B, but resized to aid visibility.
@@
Raw egocentric gaze coordinates and homography-transformed gaze coordinates were projected onto each viewer’s egoview and the shared centralview, respectively, after removal of blinks. The visualisations (presented in Figure 5) enable data validation and demonstrate accurate projective transformation using the SuperGlue features identified in the Homography module (see 3.2.6). The transforms computed using our SuperGlue + RANSAC approach handle video artefacts, such as motion blur (from sudden head movements), poor lighting/texture (frames with large dark regions, shadows, lens flare, glare etc.), obstructions (audience members in previous rows), and scale variations (different resolution and zoom in centralview and egoviews) reliably. However, further optimisations and investigations would be needed to identify edge cases in the homography estimation (we discuss these optimisations in 5.1.1). Figure 5 displays still, single frame examples of our transformed data; an example video can be viewed here: https://tinyurl.com/multipersonET. In the first part of this video example (tableau: “canjo”), the two performers are seated very near to each other. Even at this close distance, we can clearly distinguish audience-gaze on one or the other performer, vs. their instruments, and follow shifts in audience attention between these elements. When this “canjo” tableau ends, the two performers spread out and we see the gaze follow them around the stage. Once one of the performers starts speaking (tableau: “pod rattle incident”); the audience mostly maintain their focus on him, rarely glancing toward the other performer who is gently striking metallic objects. This video highlights the observation that the transformed gaze reliably follows the two musicians around the stage and provides proof-of-concept that our homography approach successfully projects gaze from a large audience in challenging scenes.
We analyse the 2D distribution of gaze with the help of heatmaps (see Figures 6 and 7). Figure 6 shows the raw egocentric gaze data recorded from the eye-tracking glasses in the egoview coordinate space. This egocentric gaze data is relative to the wearer’s head movements and exhibits low variance in both horizontal and vertical directions with a strong central bias. This effect could be explained by the high coupling between eye and head movements commonly observed in real-world exploration (Einhäuser et al., 2007; Tatler, 2007).
@@ 4.2.3. Analysis
Further comparisons between individual heatmaps can be made with the help of metrics from our Analysis module (see 3.2.8). Pairwise SIM and CC scores were calculated for each participant pair for each of the four event sessions. For raw egocentric gaze, the mean SIM score during film viewing was 0.35 (SD = 0.21) for G1F and 0.38 (SD = 0.19) for G2F, while the mean SIM during the performance was 0.40 (SD = 0.18) for G1P and 0.43 (SD = 0.17) for G2P. The mean CC was 0.37 (SD = 0.30) for G1F, 0.42 (SD = 0.28) for G2P, 0.43 (SD = 0.26) for G1P and 0.46 (SD = 0.25) for G2P. When using homography-transformed gaze, the scores in all sessions, for both of these metrics, increase drastically (see Figure 9; left column), nearly doubling in most cases. A 2D chart of seat positions (row, column) for each participant (see Figure 6 and 7 for audience seating positions on each day) was used to calculate euclidean distances (DI) between each participant pair. Calculated SIM scores on the homography-transformed gaze were found to be negatively correlated with the euclidean distances during the performance session on both days (i.e., the closer two participants were to each other, the stronger was the correlation between their gaze patterns and vice versa). The Pearson correlation coefficient between SIM and DI was -0.10 (p = 0.039) for G1P and -0.17 (p ¡ 0.001) for G2P.
@@
Figure 9. Gaze entropy (top left), average gaze velocity (bottom left), pairwise heatmap similarity (top right) and pairwise heatmap correlation (bottom right) measures computed for each of the two sessions on both days. Coloured dots in the left plots represent individual participant-wise entropy and velocity measures. Coloured dots in the right plots represent SIM and CC for all unique participant-pairs. Black dots in all plots represent means and error bars represent confidence intervals for the respective mean values. Gaze entropy and velocity give a measure of individual gaze variability for each participant and heatmap similarity (SIM) and correlation (CC) provide an estimate of similarities within participants’ gaze exploration for each session.
@@ 5. General Discussion
The proposed method conceptualised and tackled key challenges related to multi-person eye-tracking, yet a major external dependency of our framework is the choice of eye-tracking hardware. We acknowledge that the availability/development of required functionality in existing eye trackers and affordability are still substantial barriers. We chose Pupil Labs Neon devices in our utility test over market competitors because of their remote API control, minimal calibration requirements, and the availability of open-source gaze-processing components. However, in our initial tests we encountered an array of operational issues on the proprietary vendor application that severely hindered reliable recording from the devices. Solving these issues required hours of testing, troubleshooting or downscaling. For example, the Neon devices are also capable of recording device orientation and movement at 110 Hz using the onboard 9-DoF Inertial Measurement Unit (IMU), as well as stereo audio from on-board microphones; we had to omit these data streams to allow reliable recording of the gaze and video data streams, which were our priority. IMU sensor failures would crash the mobile application, even though the eye data were intact, and audio recording could not be reliably started remotely through API calls, as it requires the android device to be unlocked. While these issues are actively being resolved with the manufacturer’s assistance, they point out the very incipient stage of multi-person eye-tracking hardware/software, which are, traditionally, only tested for single person use cases. Moving forward, the technology would benefit from rigorous testing at the scale we propose in this paper. Further, the availability of open-source implementations would allow higher customizability for application specific requirements and thereby contribute to its adoption in novel settings.
5.1. Limitations and Future Directions
5.1.1. Edge-cases in gaze projection
Future objectives for our framework would be to further investigate and optimise gaze mapping accuracy in different environments. In our previous pilot study (Saxena and Fink, 2023), we identified a few edge case scenarios that pose a challenge to the homography transformations. In summary, the SuperGlue + RANSAC approach demonstrated robustness to head movements, illumination changes, and scene occlusions, as observed in our current utility study as well. However, close distances between the observer and scene as well as duplicate objects in a scene can result in false predictions. Systematically identifying these edge failure cases in challenging environments and addressing them—either through software and algorithmic updates or by establishing well-defined best practices for experimental setups—could significantly improve the robustness and reliability of gaze projections. To automate the collection of ground-truth gaze locations in these investigations, fiducial markers can be used by guiding participants to fixate on specific targets––as demonstrated in Appendix A.
5.1.2. Alternatives to planar homography
The assumption of a substantial distance between the observer(s) and the scene is a theoretical limitation of the current homography-based gaze projection as it projects a plane at sufficiently large (infinite) distance from one view to another. This assumption allows for more reliable and calibration-free projections, particularly when dealing with moving cameras, such as the head-mounted egoview, and dynamic (non-static) scenes, such as a music performance. We show the validity of this assumption for our current setup in appendix A. Future optimisations can also consider multi-view homography where the scene is captured by multiple centralview cameras and each egoview camera is related with the centralview corresponding to its position. This strategy would enable more precise cross-view analyses and spatial consistency in wide-view environments. To eliminate the 2D plane assumption, two alternative approaches are possible. Given the intrinsic parameters of the two cameras (egoview and centralview) and a common 3D coordinate system, a 2D point from the egoview can be projected to the 3D world coordinate system and then to the 2D centralview. This approach relies on pre-existing information about the locations of objects in the 3D world space, which is highly unfeasible, particularly in dynamic scenes. The second approach involves calibrating extrinsic positions of each egoview-centralview camera pair and projecting the 2D point from one camera’s view to another using epipolar geometry. This triangulation approach provides a general (ideal) solution, particularly for scenes with varying depths or close viewing distances where the homography planar assumption is not valid. However, it relies on accurate calibration parameters between the two cameras which is challenging with non-stationary cameras such as the head-mounted egoview cameras. As a result, the application of this approach relies on accurate and frequent re-calibration of the egoview cameras and prominent (unobstructed, well-lighted, strategically placed) visual markers that can be detected in the dynamic scene. These requirements might often interfere with the experimental setup, since the control of lighting and marker placement is not always possible. Additionally, markers create a visual distraction for viewers that would interfere with scientific investigations of visual attention. As an alternative to visual markers, calibration with invisible infrared markers (Dogan et al., 2022) or pose estimation of the moving egoview camera using algorithms such as visualSLAM (Chen et al., 2022) and optical flow (Beauchemin and Barron, 1995) is also possible. A potentially more direct approach would be to infer the pose of the egoview camera at each timepoint by using Inertial Measurement Unit (IMU) sensors. The eye-tracking hardware used in our utility study (see 4.1.1) are equipped with on-device 9DOF IMU. However, as mentioned previously, the IMU stream had to be disabled in favour of gaze data reliability. Overall, triangulation based methods might offer benefits, such as depth compensation and runtime efficiency, warranting future investigations to evaluate their tradeoffs in performance, accuracy, and ecological validity compared to the current homography approach.
5.1.3. Optimisations for Streaming mode of operation
The current utility study evaluated the recording (offline) mode of SocialEyes (defined in 3.1) where accuracy and resolution of recorded data is prioritised over run-time efficiency. While certain components of the recording mode, like real-time device monitoring, are directly transferable to the streaming (online) mode, future optimisations for real-time gaze projection, data acquisition/transmission, and hardware-accelerated video decoding for multi-camera streaming need to be tested at a scale similar to our current test. Below, we share insights and recommendations on these optimisations based on the current testing experience.
Real-time gaze projection could be realised with the existing SuperGlue + RANSAC method or the alternative approaches discussed in 5.1.2, tuning for the desirable accuracy-performance trade-off. For instance, marker-based plane detection or feature tracking using methods like Kanade–Lucas–Tomasi (Shi et al., 1994) can offer better run-time efficiency with slightly less accuracy or reliability of projective transformation. Similarly, multi-view homography and stereo calibration based 3D triangulation can offer better accuracy and robustness in diverse environments but might be more computationally intensive.
Data storage and transmission could be optimised in recording mode using the IMU data by only retaining the frames corresponding to substantial rotational or translational changes. These motion changes can also be detected from the egoview camera signal by calculating pixelwise change or optical flow, which reduces the requirement of an additional IMU data stream. However, detecting motion from pixels as opposed to IMU sensors would also be more sensitive to video artefacts such as lighting changes. In contrast, for the recording mode used in the current study, we ensured data reliability and reduced network utilisation by storing data locally on each eye-tracking device. Nonetheless, these strategies of motion detection might still be applied to optimise offline computation time by detecting robust features only on key frames with high inter-frame motion and tracking them on the remaining frames with low motion.
Streaming video data concurrently from multiple egoview cameras is typically challenging with software decoding due to increased load on CPU. For instance, in our current eye-tracking hardware, egoviews are sent as Full HD H.264 streams over the network. Our initial tests on a server with an Intel Core i9-13900KF CPU and a NVIDIA GeForce RTX 4090 GPU resulted in latency issues and dropped key-frames when attempting more than 5 concurrent video streams decoded on CPU (default manufacturer implementation). However, hardware-accelerated decoding with NVDEC (custom implementation with FFMPEG) was able to reliably stream all 30 egoviews concurrently with 10-30% GPU utilisation. Future testing for longer durations and optimizations, such as reducing stream bitrates—which are currently constrained by the proprietary nature of the eye-tracking application—and refining the decoding methods, will be crucial to ensure concurrent stream decoding.
5.1.4. Scope enhancement
@@
Finally, the SocialEyes framework can be extended to support additional social environments and hardware vendors. The current paper demonstrates gaze mapping in scenes where multiple people share a scene view; however, the application could be scaled up to social settings where individuals share different views of the same scene using Multi-view Multi-object detection and tracking (Taj and Cavallaro, 2010). Creative applications, such as gaze-contingent interactions and sonification (Hornof, 2014), provide an exciting avenue for exploration as well. Future development and applications can also incorporate multiple sensor types, such as physiological or IOT sensors, to record and analyse synchronous multimodal measures. We look forward to applying this framework to study attention and subjective experiences in immersive, social contexts, and plan to open-source the code to motivate collaboration and application of our framework in novel contexts.
@@ 7. Funding
Appendix A Appendix: Evaluation of homography projection at varying scene depths
In the current study, gaze projection is achieved using homography, which projects the visual scene on stage as a 2D plane from the egoview to the centralview perspective. This planar mapping assumes that the scene can be approximated as a plane at a sufficiently large (effectively infinite) distance from the camera. To assess the validity of this plane-at-infinity assumption, we conducted a post-hoc validation procedure. Note that this procedure and associated results do not directly relate to the performance and design of the SocialEyes framework presented in this paper. The aim here is to quantify the impact of varying scene depths on the accuracy of the current homography-based projections, providing additional insights for interested readers.
Figure 13. A: The visual marker was was placed at 9 different locations spread accross the stage. Red ’X’s in the image mark these positions. Distances between the positions are marked in inches. The top of the image represents the back of the stage and the bottom represents the front. Subscripted numbers 1-9 represent the sequence of presentation. B: An image of the visual marker used in the study.
[…]
Problem setting
A.1. Method
A.1.1. Apparatus
The validation procedure was conducted with the existing Pupil Labs Neon eye-trackers in the same physical space as the Utility Test in 4. Centralview was captured with an ArduCam EK035 camera, featuring a global shutter and frame-wise hardware timestamping. An ArUco marker–25.5 inches wide and 21 inches tall–was mounted on a sturdy stand and displayed to the audience at 9 different positions on the stage. Neon devices were wirelessly connected to the local network and the current SocialEyes implementation was used for data collection and analysis.
A.1.2. Procedure
19 participants were recruited during a public event at LIVELab, McMaster University. Participants seated themselves randomly among the audience seats and were equipped with the eye-trackers before the event started. The seating rows began at a distance of 134 inches from the stage front, with 50 inches of gap between each consecutive row. The validation procedure was conducted at the end of the hour-long event and was completed in roughly 3 minutes. During the procedure, an experimenter serially placed the ArUco marker (Figure 13B) at each of the positions–$X_{1}$ to $X_{9}$–presented in Figure 13A. Participants were instructed to look at the ArUco target, ensuring that the target was visible in their egoviews.
A.1.3. Data Analysis
100 consecutive frames from the centralview video were selected for each of the 9 presentation locations. ArUco markers were detected in time-synchronised egoviews of each participant and transformed to centralview using SocialEyes’ Homography module (Section 3.2.6). Projection errors were calculated as the Euclidean distances between the transformed and the detected marker centers in the centralview (640x480 resolution). Data from five devices were excluded due to more than 40% missing projections, either due to failure in marker detection or due to inadequate number of features detected for homography.
A.2. Results
Errors for each presentation location and depth are presented in Tables 2 and 2 respectively. Aggregated mean projection error reduced with an increase in depth, however, pairwise comparisons using a Wilcoxon signed-rank test showed no significant differences between depth levels 1 and 2 (W = 86.00, p = 1.00), 1 and 3 (W = 39.00, p = 0.07), or 2 and 3 (W = 44.00, p = 0.13), where depth level 1 is closest to the audience and 3 is the farthest.
Position Frames Mean Std
1 882 47.83 98.94
2 1090 49.35 104.60
3 1106 32.93 39.82
4 1141 29.62 41.65
5 1248 41.87 74.26
6 1303 55.22 45.03
7 962 39.44 44.43
8 1184 30.29 33.84
9 969 35.01 32.84
Table 1. Errors calculated for each presentation location. The position indices correspond to positions in Figure 13.
Depth Frames Mean Std
1 3147 48.32 65.02
2 3522 40.29 76.04
3 3216 32.38 38.60
Table 2. Errors aggregated for each of the 3 depth levels. Depth level 1 was closest to the audience (front of stage) and 3 was the farthest (back of stage).
A.3. Discussion
This validation procedure evaluated the projection error using our proposed homography approach, at varying scene depths. A fiducial marker was recorded from multiple egoviews spread across the audience seating space and the projected marker-centre from each egoview was compared to the ground-truth marker-centre in centralview. Projection errors across three different scene depths (each comprising of varying viewing angles) revealed no significant differences suggesting that the homography planar assumption is valid for the current setup.
A limitation in the current procedure is the poor image quality of the centralview camera due to the unavailability of appropriate Image Signal Processing (ISP) techniques for the utilised sensor. We used the ArduCam camera in this procedure for its benefits of global shutter and hardware timestamping that resulted in precise time synchronisation, however, the lack of ISP configuration for the sensor resulted in added noise to the homography projections. Note that this centralview camera was different from the one used in the Utility Study reported in the main paper, which did not have any ISP issue.
In summary, despite the added sensor noise in centralview data, the homogarphy planar assumption remains valid. The projection errors did not show significant variation across different scene depths. For future applications, it is important to ensure sufficient distance between the scene and observers for the assumption to hold, as projection errors may increase significantly at shorter distances.